please help me solve this word problem: a plane flies 450 miles with the wind and 330 miles against the wind in the same lenghth of time. If the speed of the wind is 20 mph, what is the speed of the plane in still air ?
Answer by stanbon(75887)

please help me solve this word problem: a plane flies 450 miles with the wind and 330 miles against the wind in the same length of time. If the speed of the wind is 20 mph, what is the speed of the plane in still air ?
--------------------------
With wind DATA:
distance = 450 miles ; rate = p+20 mph ; time = 450/(p+20) hrs.
---
Against wind DATA:
distance = 330 miles ; rate = p-20 mph ; time = 330/(p-20) hrs.
====================
Equation:
time = time
450/(p+20) = 330/(p-20)
-----
450p-20*450 = 330p+20*330
----
120p = 780*20
p = 130 mph (speed of the plane in still air)
=====================
Cheers,
Stan H.